User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!



VSN: HOB+0096 ( +9464 GB)

  • 1700 UT: First scan missed due to slow change over from pulsar observations. (DH)
  • 1700 UT: Cryo pressures high, 20K stage at 90K. Driving to Mt P to vacuum. (DH)
  • 1800 UT: Vacuumed cryos. Temps slowly decreasing (DH).
  • 1517UT SX para pressure out of range. Reading is 4.471. 20K stage at 80.66 hopefully it doesn't get too high in the next 2 hours. (JS)
  • 1600UT Pressure up to 4.9 and 20k temp at 79K advised to leave for the 1hour left in experiment (JS)


VSN: HAY-0029 (+0 GB)

NB: still using mk5-2ke, use clkoff-2 to check delays

  • Autocorrelation has large spikes in bands 01 and 14 (Dave)


VSN: TR-00009 (+0 GB)

* The humidity monitor is not working at the moment (IJ)

You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/r1709.txt · Last modified: 2015/10/14 19:27 by Jesse Swan